OpenStudy (ja1):

Explain how you would find the exact value of:

OpenStudy (ja1):

\[\frac{ \sqrt{2} }{ \sqrt{8} }\]

OpenStudy (psymon):

\[\frac{ \sqrt{2} }{ \sqrt{8} }=\sqrt{\frac{ 2 }{ 8 }}\] Think you can find it from there? :P

OpenStudy (anonymous):

1/2

OpenStudy (ja1):

I actually prefer doing it the other way :/

OpenStudy (luigi0210):

Multiply top and bottom by rad 8

OpenStudy (anonymous):

LOL

OpenStudy (luigi0210):

Wait, is this trig?

OpenStudy (ja1):

Wouldn't you simplify 8 first?

OpenStudy (ja1):

no

OpenStudy (psymon):

\[\frac{ \sqrt{2} }{ \sqrt{8} }*\frac{ \sqrt{8} }{ \sqrt{8} }?\]

OpenStudy (ja1):

I would get: \[\frac{ \sqrt{2} }{ 2\sqrt{2} }\]

OpenStudy (ja1):

Oh wait a second.....

OpenStudy (psymon):

Which becomes 1/2.

OpenStudy (ja1):

*Facebrick*

OpenStudy (luigi0210):

:3

OpenStudy (ja1):

Wait, wouldn't the two √2 cancel out and leave 2?

OpenStudy (psymon):

\[\frac{ \sqrt{2} }{ \sqrt{8} }*\frac{ (╯°□°)╯︵ ┻━┻ }{ \infty } \]

OpenStudy (ja1):

sorry @Psymon I like doing it the other way :3

OpenStudy (psymon):

No worries. We have our preferred ways xD

OpenStudy (ja1):

Oh ok I see now, when the two √2 cancel out you put a 1 in the numerator

OpenStudy (ja1):

Thus getting 1/2

OpenStudy (psymon):

Yep.

OpenStudy (ja1):

I got it now xD

OpenStudy (ja1):

Thanks all I was braindead for a second there

OpenStudy (psymon):

Glad ya got it xD

OpenStudy (ja1):

C: